OpenStudy (acacia):

Given the hyperbola (y2/16) – (x2/9) = 1, find the vertices

OpenStudy (anonymous):

\[\frac{y^2}{4^2}-\frac{x^2}{3^2}=1\] is a start

OpenStudy (anonymous):

you need to know two things, first that the vertex is \((0,0)\) and second, that because the \(y^2\) term comes first it looks like this 

OpenStudy (anonymous):

this tells you that the vertex is 4 units (the 4 because of the \(\frac{y^2}{4^2}\) term) above and below the vertex